OpenStudy (anonymous):

*MEDAL* What is the value of x? 4x - 9 = 27 x= ____

OpenStudy (anonymous):

@Kyliejones why are you posting these questions if you already know how to do them?

OpenStudy (anonymous):

4x=27+9

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

x=36/4=9

OpenStudy (anonymous):

x=9